Please help. I'm having trouble with this problem.
The volume of a cube is given by V = s3, where s is the length of a side. Find the length of a side of a cube if the Volume is 1000 cm3.
Answer:

Answer by venugopalramana(3286)

Please help. I'm having trouble with this problem.
The volume of a cube is given by V = s3, where s is the length of a side. Find the length of a side of a cube if the Volume is 1000 cm3.
V=1000 CU.CM=10*10*10=S*S*S
HENCE S=10 CM
Answer: